ABSTRACT. For a set-valued mapping, the relationships between lower semicontinuity, almost lower semicontinuity, and the existence of various kinds of continuous selections for the mapping are explored. For the space C_{0}(T) and L_{1}, intrinsic characterizations are given of the one-dimensional subspaces whose metric projections admit continuous selections.